OpenStudy (batmaaaanxd):

[Medal & Fan] Prove the following identity: sin4A + cos4A/1-2sin2A cos2A = 1

OpenStudy (batmaaaanxd):

@mathmate

OpenStudy (batmaaaanxd):

@.Sam.

OpenStudy (mathmate):

sin4A + cos4A/1-2sin2A cos2A = 1 As is, mathematically actually means \(sin4A + \frac{cos4A}{1}-2sin2A cos2A = 1\) which is probably not what you mean. So inserting parentheses will make it crystal clear: \((sin4A + cos4A)/(1-2sin2A cos2A) = 1\) or even better, use LaTeX \(\Large \frac{\sin(4A) + \cos(4A)}{1-2\sin(2A) \cos(2A)} = 1 \) You will notice that in LaTeX and in typeset documents, parentheses are not shown, BUT when transcribing to a single line, the parentheses are necessary. Having said all that, the above cannot be proved, because it is not an identity. The value on the left-hand side varies with the angle A. When A equals 1 radian, the left-hand side evaluates to -0.8028, i.e. does not equal 1. I suggest you double-check all mathematical expressions/equations before posting.

OpenStudy (batmaaaanxd):

@mathmate Thanks for your response! And sorry, my bad. :/

OpenStudy (mathmate):

hint: try sum and difference formulas product formulas The following link is full of treasures for what you need to solve the problem: http://www.regentsprep.org/regents/math/algtrig/att14/formulalesson.htm

OpenStudy (batmaaaanxd):

@mathmate Thanks! I'll go through the site

OpenStudy (mathmate):

@batmaaaanxD no problem!